find two consecutive odd integers such that 5 times the first integer is 12 more than 3 times the second.
Answer by stanbon(75887)

find two consecutive odd integers such that 5 times the first integer is 12 more than 3 times the second.
----------------------------
Let the consecutive odd integers be x and x+2.
EQUATION:
5x=3(x+2)+12
5x = 3x+18
2x = 18
x=9 (1st odd integer)
x+2 = 11 (2nd odd integer)
==================
Cheers,
Stan H.